User name B. Craver
Log entry time 00:07:14 on August11,2003
Entry number 110060
keyword=8/10 Swing shift summary
16:00 Begin shift waiting for beam.
17:30 Request Controlled Access
17:45 Hall in Controlled Access, Alexandre enters to check collimators
18:00 We notice that the N2 level for the Q3 has dropped to 0.
18:10 We page Scot Spiegel about the Q3.
18:20 Still no answer from Scot, but N2 level rises back to normal. The stripchart
shows that the loss of pressure has been occurring periodically for the last
18:40 Alexandre exits hall and we request return to Beam Permit.